# Machine Learning-Based Ultrasonic Weld Characterization Using Hierarchical Wave Modeling and Diffusion-Driven Distribution Alignment

### v1

@@ Appendix A Weldline Parameterization Functions
The bead function is selected to provide realistic impedance profiles around the weld-line. It’s upon the path $\Gamma_{c}$ and radius function $f_{r}(d_{\perp})$, where $d_{\perp}$ is the perpinduclar distance from the weld path,
@@
where $R$ is the weld radius and $r_{f}-\alpha R$ a small fillet radius to ensure smoothness in the domain. Moreover, real welds have more appreciable imepdenence mismatch laong weldline boundaries than those modeled by (38), and we therefore consider an additional boundary wavespeed reduction term as,
@@ B.1   Elastodynamic Equations
The elastodynamic model of Eq (1) was solved on via a Galerkin-FE approximation. Let $V_{h}\subset V:=\{\bm{v}\in[H^{1}(\Omega)]^{3}|\bm{v}=0\ \text{on}\ \Gamma_{\mathsf{D}}\}$ be the vector-valued quadratic ($P_{2}$) Lagrange space with homogeneous Dirichlet conditions on $\Gamma_{\mathsf{D}}$. We seek a trial function $\bm{u}_{h}\in V_{h}$ such that for all $\bm{v}_{h}\in V_{h}$,
@@ B.2   Helmholtz Equation
Let $W_{h}\subset W=\{v\in H^{1}(\Omega^{\mathsf{R}}):v=0\text{ on }\Gamma_{\mathsf{D}}^{\mathsf{R}}\}$ be the scalar quadratic ($P_{2}$) Lagrange space with homogeneous conditions on $\Gamma_{\mathsf{D}}^{\mathsf{R}}$. We seek $\psi_{h}^{(m,n)}\in W_{h}$ such that or all $v_{h}\in W_{h}$
@@ Appendix C Comparison of Conditioned DDPM to DnCNN
Figure 15: Performance on ML-processed experimental measurements using (a,b) DDPM conditional generation and (c,d) DnCNN denoising, with with (a,c) no additive white Gaussian noise (AWGN) and (b,d) 40% AWGN.

### v2

@@ Appendix A Weldline Parameterization Functions
The bead function is selected to provide realistic impedance profiles around the weld-line. It’s upon the path $\Gamma_{c}$ and radius function $f_{r}(d_{\perp})$, where $d_{\perp}$ is the perpendicular distance from the weld path,
@@
where $R$ is the weld radius and $r_{f}-\alpha R$ a small fillet radius to ensure smoothness in the domain. Moreover, real welds have more appreciable impedance mismatch along weldline boundaries than those modeled by (38), and we therefore consider an additional boundary wavespeed reduction term as,
@@ B.1   Elastodynamic Equations
The elastodynamic model of Eq (1) was solved via a Galerkin-FE approximation. Let $V_{h}\subset V:=\{\bm{v}\in[H^{1}(\Omega)]^{3}|\bm{v}=0\ \text{on}\ \Gamma_{\mathsf{D}}\}$ be the vector-valued quadratic ($P_{2}$) Lagrange space with homogeneous Dirichlet conditions on $\Gamma_{\mathsf{D}}$. We seek a trial function $\bm{u}_{h}\in V_{h}$ such that for all $\bm{v}_{h}\in V_{h}$,
@@ B.2   Helmholtz Equation
Let $W_{h}\subset W=\{v\in H^{1}(\Omega^{\mathsf{R}}):v=0\text{ on }\Gamma_{\mathsf{D}}^{\mathsf{R}}\}$ be the scalar quadratic ($P_{2}$) Lagrange space with homogeneous conditions on $\Gamma_{\mathsf{D}}^{\mathsf{R}}$. We seek $\psi_{h}^{(m,n)}\in W_{h}$ such that for all $v_{h}\in W_{h}$
@@ Appendix C Comparison of Conditioned DDPM to DnCNN
Figure 15: Performance on ML-processed experimental measurements using (a,b) DDPM conditional generation and (c,d) DnCNN denoising, with (a,c) no additive white Gaussian noise (AWGN) and (b,d) 40% AWGN.